OpenStudy (anonymous):

Give reasons for each step of the deductive argument that if x – 5 = 2(x + 6), then x = -17. x – 5 = 2(x + 6) Original problem x – 5 = 2x + 12 1) -5 = x + 12 2) -17 = x 3) x = -17 4) Answer choices: Reflexive property Symmetric property Transitive property Addition property of equality Subtraction property of equality Multiplication property of equality Division property of equality Substitution property Distributive property

OpenStudy (texaschic101):

x - 5 = 2(x + 6) --- original problem x - 5 = 2x + 12 --- distributive property -5 = x + 12 -- result of subtracting x ...subtraction property of equality -17 = x -- result of subtracting 12...subtraction property x = -17 -- result of multiplying by -1 ...multiplication property......or I guess it could be reflexive property...not 100% sure.

OpenStudy (anonymous):

Thanks, that was helpful!